# Physical conditions that lead to the detection of the pair annihilation line in the BOAT GRB221009A

### v1

@@ 1 Introduction
Although the creation of a large number of pairs was predicted a long time ago (Pe’er & Waxman, 2004; Pe’er et al., 2006; Ioka et al., 2007), a firm observational confirmation of the existence of an expected pair annihilation line in the observed GRB spectra was so far lacking. This fact can partially be explained as that the blue-shifted line is expected to be detected in the hundred’s of MeVs (the Fermi-LAT) range, in which the detector’s sensitivity is poorer than at lower energy bands. In addition, due to relativistic motion, this line may be smeared if the Lorentz factor in different jet regions is somewhat different. Nonetheless, the lack of detection of this line seem to contradict with theoretical expectations.
@@ 2 Key observations and basic model assumptions
The observed hard X-ray lightcurve of GRB22109A is very long, lasting for several hundreds of seconds. It is variable, with variability time of a few seconds, and show multiple peaks. The highest peak occurs at $\sim 244$ s after the GRB trigger. In fact, this peak was so bright that the Fermi-GBM detector was saturated in between the time interval 219-277 s (Edvige Ravasio et al., 2023). The physical origin of this peak is uncertain. In particular, the radius in which energy is dissipated is unknown. However, the fact that the observed spectra is broad, and extends to high energies suggests that the energy dissipation that produces this peak occurs above the photosphere. A likely possibility is that a narrow Poynting flux dominated jet dissipate energy at an emission radius beyond $10^{16}$ cm (see Zhang et al. (2024a) for arguments).
@@ 3.3 High optical depth for pair production
${n^{\prime}_{e^{\pm}}}^{2}={L\over(4\pi r^{2}\Gamma^{2}Am_{e}c^{3}t^{\prime}_{dyn})}{1\over c\sigma_{T}}={L\sigma_{T}t_{dyn}\over(4\pi r^{2}\Gamma^{2}Am_{e}c^{2})}\left({1\over c\sigma_{T}t^{\prime}_{dyn}}\right)^{2}={\tau_{\gamma\gamma}A^{2}\over(c\sigma_{T}t^{\prime}_{dyn})^{2}},$ (21)
@@
$n^{\prime}_{e^{\pm}}\sim{\tau_{\gamma\gamma}^{1/2}A\over c\sigma_{T}t^{\prime}_{dyn}}=2.5\times 10^{9}~{}A^{3/2}~{}L_{54}^{1/2}~{}\Delta t_{1}^{-3/2}~{}\left({\Gamma\over 600}\right)^{-7/2}~{}{\rm cm^{-3}}.$ (22)
@@
$\tau_{e^{\pm}}=\Delta rn^{\prime}_{e^{\pm}}\sigma_{T}=\tau_{\gamma\gamma}^{1/2}A,$ (23)
namely, at any given radius, $\tau_{\gamma\gamma}$ and $\tau_{e^{\pm}}$ are either both greater or smaller than unity.
@@
Figure 3: Rate of pair annihilation as a function of the Lorentz factor, $\Gamma$, from Equations 19 and 24
At low $\Gamma$ the system is in the high optical depth to pair production regime, and the rate of annihilation is proportional to $\Gamma^{-2}$, while at higher values of $\Gamma$ (the low optical depth regime), it sharply decays as $\Gamma^{-17}$. Solid lines are for $A=1$, while dashed lines are for $A=8$ (note that $A=10$ would result in lines overlap).
@@ 4.1 Low optical depth for pair production
The pair annihilation rate calculated in Equation (26) is to be compared to the theoretical expectation given in Equation (19). Basic algebra reveals that the two results match for $\Gamma=620~{}L_{54}^{2/7}$ and $\Delta t=3.0$ s. Assuming these value of the Lorentz factor, the dissipation radius is $r=3.5\times 10^{16}$ cm, which is much higher than the photospheric radius, at $r_{ph}=2.4\times 10^{12}$ cm. The optical depth for pair production is $\tau_{\gamma\gamma}=0.26~{}A$, i.e., close to unity, and for pair annihilation it is $\tau_{e^{\pm}}=0.068$. For these parameters, The time to reach steady state, $t^{\prime}_{ss}=7.1\times 10^{3}$ s is indeed longer than the dynamical time, $t^{\prime}_{dyn}=1.86\times 10^{3}$ s (although by a factor of a few only), implying that the approximation we used above is valid.
@@ Acknowledgements
AP acknowledge the support from the European Union (EU) via ERC consolidator grant 773062 (O.M.J.).

### v2

@@ 1 Introduction
Although the creation of a large number of pairs was predicted a long time ago (Pe’er & Waxman, 2004; Pe’er et al., 2006; Ioka et al., 2007; Murase & Ioka, 2008), a firm observational confirmation of the existence of an expected pair annihilation line in the observed GRB spectra was so far lacking. This fact can partially be explained as that the blue-shifted line is expected to be detected in the hundred’s of MeVs (the Fermi-LAT) range, in which the detector’s sensitivity is poorer than at lower energy bands. In addition, due to relativistic motion, this line may be smeared if the Lorentz factor in different jet regions is somewhat different. Nonetheless, the lack of detection of this line seem to contradict with theoretical expectations.
@@ 2 Key observations and basic model assumptions
The observed hard X-ray lightcurve of GRB22109A is very long, lasting for several hundreds of seconds. It is variable, with variability time of a few seconds, and show multiple peaks. The highest peak occurs at $\sim 244$ s after the GRB trigger. In fact, this peak was so bright that the Fermi-GBM detector was saturated in between the time interval 219-277 s (Edvige Ravasio et al., 2023). The physical origin of this peak is uncertain. In particular, the radius in which energy is dissipated is unknown. However, the fact that the observed spectra is broad, and extends to high energies suggests that the energy dissipation that produces this peak occurs above the photosphere. A likely possibility is that a narrow Poynting flux dominated jet dissipate energy at an emission radius beyond $10^{16}$ cm (see Dai et al. (2023); Zhang et al. (2024a) for arguments) [^1] The broad-band afterglow emission requires the existence of two jet components, with a narrow core and a structured wing (see Zhang et al. (2024a) for the arguments supporting such a jet structured and detailed modeling by Sato et al. (2023); Zheng et al. (2024); Ren et al. (2024); Zhang et al. (2023)). The prompt emission is likely related to energy dissipation within the narrow jet.
@@ 3.3 High optical depth for pair production
${n^{\prime}_{e^{\pm}}}^{2}={L\over(4\pi r^{2}\Gamma^{2}Am_{e}c^{3}t^{\prime}_{dyn})}{1\over c\sigma_{T}}={L\sigma_{T}t^{\prime}_{dyn}\over(4\pi r^{2}\Gamma^{2}Am_{e}c^{2})}\left({1\over c\sigma_{T}t^{\prime}_{dyn}}\right)^{2}={\tau_{\gamma\gamma}\over A^{2}(c\sigma_{T}t^{\prime}_{dyn})^{2}},$ (21)
@@
$n^{\prime}_{e^{\pm}}\sim{\tau_{\gamma\gamma}^{1/2}\over Ac\sigma_{T}t^{\prime}_{dyn}}=2.5\times 10^{9}~{}A^{-1/2}~{}L_{54}^{1/2}~{}\Delta t_{1}^{-3/2}~{}\left({\Gamma\over 600}\right)^{-7/2}~{}{\rm cm^{-3}}.$ (22)
@@
$\tau_{e^{\pm}}=\Delta rn^{\prime}_{e^{\pm}}\sigma_{T}=\tau_{\gamma\gamma}^{1/2}A^{-1},$ (23)
namely, at any given radius, $\tau_{\gamma\gamma}$ and $\tau_{e^{\pm}}$ are either both greater or smaller than unity (for $A=1$).
@@
Figure 3: Rate of pair annihilation as a function of the Lorentz factor, $\Gamma$, from Equations 19 and 24. At low $\Gamma$ the system is in the high optical depth to pair production regime, and the rate of annihilation is proportional to $\Gamma^{-2}$, while at higher values of $\Gamma$ (the low optical depth regime), it sharply decays as $\Gamma^{-17}$. Solid lines are for $A=1$, while dashed lines are for $A=8$ (note that $A=10$ would result in lines overlap).
@@ 4.1 Low optical depth for pair production
The pair annihilation rate calculated in Equation (26) is to be compared to the theoretical expectation given in Equation (19). Basic algebra reveals that the two results match for $\Gamma=620~{}L_{54}^{2/7}$ and $\Delta t=3.0$ s. Assuming these value of the Lorentz factor, the dissipation radius is $r=3.5\times 10^{16}$ cm, which is much higher than the photospheric radius, at $r_{ph}=2.4\times 10^{12}$ cm. The optical depth for pair production is $\tau_{\gamma\gamma}=0.26~{}A$, i.e., close to unity, and for pair annihilation it is $\tau_{e^{\pm}}=0.068~{}A^{2}$. For these parameters, The time to reach steady state, $t^{\prime}_{ss}=7.1\times 10^{3}$ s is indeed longer than the dynamical time, $t^{\prime}_{dyn}=1.86\times 10^{3}$ s (although by a factor of a few only), implying that the approximation we used above is valid.
@@ 4.2 High optical depth for pair production
The cross section for photon scattering by the created pairs in the plasma, $\sigma_{T}$, is comparable to the cross section of annihilation of cold pair. Therefore, Equation 23 provides a very good proxy for the optical depth for scattering by the created pairs. Using Equation 6 and the observational constraint set above, the optical depth for photon scattering is given by
$\tau_{sc.}=0.55~{}A^{-1/2}~{}L_{54}^{1/2}~{}\left({\Gamma\over 600}\right)^{-2}<0.12~{}L_{54}^{1/3}~{}\left({\Gamma\over 600}\right)^{-2},$ (29)
where in the last equality we used equation 27. One therefore finds that for most of parameter space regime of interest, this optical depth is smaller than unity, and even for the most extreme cases of $L=10^{54}$ erg/s and $\Gamma\gtrsim 100$, it is not more than a few at most. (Lower values of Lorentz factor will imply that the emission radius is close to the photospheric radius, and the resulting spectra will become quasi-thermal. The observed spectra therefore exclude this parameter space). One therefore concludes that scattering by the pairs is not expected to have a major effect on the observed line properties, such as its observed brightness or width.
@@ Acknowledgements
AP acknowledges the support from the European Union (EU) via ERC consolidator grant 773062 (O.M.J.). BZ acknowledges NASA 80NSSC23M0104 and the Nevada Center for Astrophysics for support.